# MIMO: Controllable Character Video Synthesis with Spatial Decomposed Modeling

### v1

@@ 3 Experimental Results
Dataset. We create a human video dataset called HUD-7K to train our model. This dataset consists of $5K$ real character videos and $2K$ synthetic character animations. The former does not require any annotations and can be automatically decomposed to various spatial attributes via our scheme. To enlarge the range of the real dataset, we also synthesize $2K$ videos by rendering character animations in complex motions under multiple camera views, utilizing En3D [21]. These synthetic videos are equipped with accurate annotations due to completely controlled production.
Figure 5: Results of animating diverse characters from a single reference image.
Figure 6: Results of synthesizing avatar animations with novel 3D motions, which are retrieved from the motion database or extracted from in-the-wild human videos.
Figure 7: Results of synthesizing avatar animations with interactive scenes, which are extracted from in-the-wild videos.
@@ 3.1 Controllable character video synthesis
Given the target attributes of character, motion and scene, our method can generate realistic video results with their latent codes combined for guided synthesis. The target attributes can be provided by simple user inputs (e.g., single images/videos for character/scene, pose sequences from large database [20, 1] for motion) or flexibly extracted from the real-world videos, involving complicated scenes of occluded object interactions and extreme articulated motions. In the following, MIMO demonstrates that it can simultaneously achieve advanced scalability to arbitrary characters, generality to novel 3D motions, and applicability to in-the-wild scenes in a unified framework.
3.1.1 Arbitrary character control
@@
As shown in Figure 5, our method can animate arbitrary characters, including realistic humans, cartoon characters and personified ones. Various body shapes of characters can be faithfully preserved due to the decoupled pose and shape parameters in our structured motion representation.
3.1.2 Novel 3D motion control
@@
To verify the generality to novel 3D motions, we test MIMO using challenging out-of-distribution pose sequences from the AMASS [20] and Mixamo [1] database, including dancing, playing and climbing (Figure 6 (a)). We also try complex spatial motions in 3D space by extracting them from in-the-wild human videos (Figure  6 (b)). Our method exhibits high robustness for these novel 3D motions under different viewpoints.
3.1.3 Interactive scene control
@@
We validate the applicability of our model to complicated real scenes by extracting scene and motion attributes from in-the-wild videos for character animation (i.e., the task of video character replacement). As shown in Figure 7, the character can be seamlessly inserted to the real scenes with natural object interactions, owing to our spatial-aware synthesis for hierarchical layers.
@@ 4 Conclusions
In this paper, we presented MIMO, a novel framework for controllable character video synthesis, which allows for flexible user control with simple attribute inputs. Our method introduces a new generative architecture which decomposes the video clip to various spatial components, and embeds their latent codes as the condition of decoder to reconstruct the video clip. Experimental results demonstrated that our method enables not only flexible character, motion and scene control, but also advanced scalability to arbitrary characters, generality to novel 3D motions, and applicability to interactive scenes. We also believed that our solution, which considers inherent 3D nature and automatically encodes the 2D video to hierarchical spatial components could inspire future researches for 3D-aware video synthesis. Furthermore, our framework is not only well suited to generate character videos but also can be potentially adapted to other controllable video synthesis tasks.

### v2

@@ 4 Experimental Results
Dataset. We create a human video dataset called HUD-7K to train our model. This dataset consists of $5K$ real character videos and $2K$ synthetic character animations. The former does not require any annotations and can be automatically decomposed to various spatial attributes via our scheme. To enlarge the range of the real dataset, we also synthesize $2K$ videos by rendering character animations in complex motions under multiple camera views, utilizing En3D [29]. These synthetic videos are equipped with accurate annotations due to completely controlled production. For the evaluation, we collect 100 in-the-wild human videos covering diverse contents (e.g., dancing, sports and movie) and randomly truncate them to 150-frame clips as test set.
Metrics. We follow [12] to evaluate our method using four standard metrics: Peak Signal-to-Noise Ratio (PSNR) [11], Structural Similarity Index Measure (SSIM) [40], Learned Perceptual Image Patch Similarity (LPIPS) [48] for image-level quality, and Fréchet Video Distance (FVD) [37] for video-level evaluation.
@@ 4.1 Controllable character video synthesis
Given the target attributes of character, motion and scene, our method can generate realistic video results with their latent codes combined for guided synthesis. The target attributes can be provided by simple user inputs (e.g., single images/videos for character/scene, pose sequences from large database [28, 1] for motion) or flexibly extracted from the real-world videos, involving complicated scenes of occluded object interactions and extreme articulated motions. In the following, MIMO demonstrates that it can simultaneously achieve advanced scalability to arbitrary characters, generality to novel 3D motions, and applicability to in-the-wild scenes in a unified framework. More results can be found in the supplemental materials (Supp).
@@
Arbitrary character control. As shown in Figure 5, our method can animate arbitrary characters, including realistic humans, cartoon characters and personified ones. Various body shapes of characters can be faithfully preserved due to the decoupled pose and shape parameters in structured motion representation.
@@
Novel 3D motion control. To verify the generality to novel 3D motions, we test MIMO using challenging out-of-distribution pose sequences from the AMASS [28] and Mixamo [1] database, including dancing, playing and climbing (Figure 5 (a)). We also try complex spatial motions by extracting them from in-the-wild videos for driving (Figure  5 (b, c)). Our method exhibits high robustness for these novel 3D motions under different viewpoints.
@@
Interactive scene control. We validate the applicability of our model to complicated real scenes by extracting both scene and motion attributes from in-the-wild videos for character animation, as a brand-new task of video character replacement. Figure 5 (c) shows that the characters can be seamlessly inserted to the real scenes with natural human-object interactions.
4.2 Comparison with state-of-the-arts
Qualitative comparison. In Figure 6, we compare the synthesis results of our method with three state-of-the-art character animation methods: Animate Anyone [12], Mimic-Motion [50] and Champ [53]. All the results of these methods are obtained by using the source codes and trained models released by authors or popular re-implements, following by fine-tuning in our training dataset. As we can see, all previous methods fail to produce extreme articulated human motions with exaggerated deformations and frequent self-occlusions (Figure 6 (a)). They also cannot handle complicated scenes of object interaction (Figure 6 (b)) and large camera movement (Figure 6 (c)). In contrast, our method tackles these challenges and gives more realistic results in both global structures and detailed textures. More video results can be found in Supp. Furthermore, our method shows its superiority that it enables directly inferring animatable avatars in free-viewpoint with inter-frame consistency to some extent, which are comparable to the results of SOTA training-based 3D method, as presented in Supp.
Quantitative comparison. Table 1 shows the comparison of our method with [12, 50, 53] in terms of the PSNR, SSIM, LPIPS and FVD metrics, respectively. Due to the presence of a certain quantity of complex cases (e.g., including spatial motions, scene interactions, and camera movements) in the test set, it’s extremely challenging to model the intricate interplay of real-world scenarios. Even so, our method demonstrated the best performance in these metrics among all methods. It outperforms previous works by a margin of at least 4.16 in terms of PSNR and 0.152 in terms of SSIM, etc. In contrast to directly learning the entire 2D video frame with only inadequate human pose annotations, MIMO decompose 2D frames into hierarchically spatial components with more expressive 3D representations. The results indicate that our method better simulates video observations of the real physical world. Considering insufficient scene modeling of previous methods, we also provide additional quantitative comparison by removing background and object for only character synthesis in Supp.
Table 1: Quantitative comparison with state-of-the-art methods in terms of PSNR, SSIM, LPIPS and FVD.
[…]
Animate Anyone [12] 21.003 0.722 0.264 304.3
Mimic-Motion [50] 20.688 0.731 0.343 289.2
Champ [53] 21.044 0.724 0.312 412.5
Ours 25.210 0.883 0.125 221.4
Figure 7: Effects of spatial decomposed modeling.
4.3 Ablation study
Spatial decomposed modeling. We assess the impact of this design by training a model via randomly selecting one frame from videos as the appearance reference without decomposed layers (w/o SDM). In this way, it fails to produce faithful background and interactive foreground, easily suffering from unstable texture distortions for large camera movements (Figure 7 (a, b)). Essentially, this instability stems from the absent guidance of scene generation, relying only on weak correlation between the scene and character movement revealed by data distribution. We also attempt to model the human and mixed scene without independent occlusion encoding (w/o occ.), and Figure 7 (c, d) shows that this variation cannot synthesize reasonable occluded objects for scene interaction without the ability to comprehend spatial layers. Figure 7 (d) and Table 2 also indicates that this decomposed strategy yield more realistic results with facial details and clothing wrinkles.
Table 2: Results of models trained by removing specific modules or replacing with alternative designs for ablation study.
[…]
Ours-w/o SDM 22.148 0.762 0.231 268.5
Ours-w/ 2D skeleton 24.326 0.842 0.186 237.2
Ours-w/ 3D maps 24.402 0.844 0.203 278.1
Ours-w/o CA 24.918 0.871 0.148 223.1
Ours 25.210 0.883 0.125 221.4
Figure 8: Effects of structured motion representation and canonical appearance transfer.
Structured motion code. To verify the effectiveness of the proposed structured motion representation (SMR), we evaluate the performance of several variants of our method by employing alternative motion formats: commonly used 2D skeleton in [12] and 3D maps in [53]. As shown in Figure 8 and Table 2, 2D skeleton ignores the occlusion relationship in bones and muscles, resulting in ambiguity for spatial motions. The 3D maps, consisting of normal map, depth map, etc., improve the pose representation capability, but still struggle for highly complex spatial motions due to undefined labels of dense body parts. Our SMR, inserting identifiable codes into structured body surfaces and projecting for 2D correspondence, provides strong articulation ability of motion in spatial space and significantly improves the model’s generalizability to novel 3D motions.
Canonical appearance transfer. This design (CA) further disentangles motion and identity in consecutive video frames with high posture correlation. It leads to more effective learning of SMR and obviously alleviates the issue of synthesis confusion between hands and feet (Figure 8 (d)).
@@ 5 Conclusions
In this paper, we presented MIMO, a novel framework for controllable character video synthesis, which allows for flexible user control with simple attribute inputs. Our method introduces a new generative architecture which decomposes the 2D video to various spatial components, and embeds their latent codes as the condition of decoder to reconstruct the video. Experimental results demonstrated that our method enables not only flexible character, motion and scene control, but also advanced scalability to arbitrary characters, generality to novel 3D motions, and applicability to interactive scenes. We also believed that our solution, which considers inherent 3D nature of video occurrence and automatically encodes the 2D video to hierarchical spatial components could inspire future researches for 3D-aware video modeling. Furthermore, our framework is not only well suited to generate character videos but also can be potentially adapted to common video synthesis tasks.